In how many ways you can select a captain and a vice captain from eleven players?

Grade:12th Pass

1 Answers

Arun
5 years ago
Dear student
 
we can select a captain in 11C1 ways and vicecaptain in 10C1
 
hence total number of ways = 11* 10 = 110
 
Method 2
 
pick two players in 11C2 ways now for captain there are 2 ways and after selecting yhere is only only one way to select vice captain.
 
hence number of ways = 11c2 * 2 = 55* 2 = 110